Alyeski Tours operates day tours of coastal glaciers in Alaska on its tour boat the Blue Glacier. Management has identified two cost drivers—the number of cruises and the number of passengers—that it uses in its budgeting and performance reports. The company publishes a schedule of day cruises that it may supplement with special sailings if there is sufficient demand. Up to 80 passengers can be accommodated on the tour boat. Data concerning the company’s cost formulas appear below:

Fixed Cost per Month Cost per Cruise Cost per Passenger
Vessel operating costs $ 5,200 $ 480.00 $ 2.00
Advertising $ 1,700
Administrative costs $ 4,300 $ 24.00 $ 1.00
Insurance $ 2,900

For example, vessel operating costs should be $5,200 per month plus $480 per cruise plus $2 per passenger. The company’s sales should average $25 per passenger. In July, the company provided 24 cruises for a total of 1,400 passengers.

Required:

Prepare the company’s flexible budget for July.
Answer #1
Flexible budget
Actual cruises 24
Actual passengers 1400
Revenue 35000 =1400*25
Expenses:
Vessel operating costs 19520 =5200+(480*24)+(1400*2)
Advertising 1700
Administrative costs 6276 =4300+(24*24)+(1400*1)
Insurance 2900
Total expenses 30396
Net Operating income 4604